3. If tan 2A = cot (A – 18°), where 2A is an acute angle, find the value of A.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
tan 2A = cot (A- 18°) We know that tan 2A = cot (90° – 2A) Substitute the above equation in the given problem ⇒ cot (90° – 2A) = cot (A -18°) Now, equate the angles, ⇒ 90° – 2A = A- 18° ⇒ 108° = 3A A = 108° / 3 Therefore, the value of A = 36°